Proceedings of the

6th Nordic Workshop on Programming Theory (Aarhus, Denmark, 17–19 October, 1994)

December 1994

This document is also available as PostScript, DVI, Text.

References

[Hen] Abstract, PostScript, DVI.

Matthew Hennessy. Higher-order processes and their models. In 6th NWPT, page 1. vi+483 pp.

[Ste] Abstract, PostScript.

Bernhard Steffen. Finite model checking and beyond. In 6th NWPT, pages 2–17. vi+483 pp.

[NS] Abstract, PostScript, DVI.

Dave S. Neilson and Ib Holm Sørensen. The B-Technologies: A system for computer aided programming. In 6th NWPT, pages 18–35. vi+483 pp.

[AHS] Abstract, PostScript, DVI.

Hosein Askari, Ole I. Hougaard, and Michael I. Schwartzbach. A framework for ad-hoc type inference. In 6th NWPT, pages 36–50. Extended Abstract.

[AI] Abstract, PostScript, DVI.

Luca Aceto and Anna Ingólfsdóttir. CPO models for a class of GSOS languages. In 6th NWPT, pages 51–66. vi+483 pp.

[AK] Abstract, PostScript.

Parosh Aziz Abdullah and Mats Kindahl. On decidability of simulation and bisimulation for lossy channel systems. In 6th NWPT, pages 67–77. vi+483 pp.

[AM] Abstract, PostScript, DVI, Relevant papers. Henrik Reif Andersen and Michael Mendler. PMC: A process algebra for real-time systems. In 6th NWPT, pages 78–79. vi+483 pp.

[ANN] Abstract, PostScript, DVI.

Torben Amtoft, Flemming Nielson, and Hanne Riis Nielson. Type and behaviour reconstruction for higher-order concurrent programs. In 6th NWPT, pages 80–95. vi+483 pp.

[BB] Abstract, PostScript, DVI. Ralph-Johan Back and Michael J. Butler. Applications of summation and product operators in the refinement calculus. In 6th NWPT, pages 96–111. vi+483 pp.

[CH] Abstract, PostScript, DVI. Vytautas Čyras and Magne Haveraaen. Programming with data dependencies: a comparison of two approaches. In 6th NWPT, pages 112–126. Some proofs are omitted due to lack of space.

[Cer] Abstract, PostScript, DVI. Kārlis Čerāns. A calculus of timed refinement. In 6th NWPT, pages 127–141. vi+483 pp.

[Che] Abstract, PostScript. Allan Cheng. Petri nets, traces, and local model checking. In 6th NWPT, pages 142–156. vi+483 pp.

[Coq] Abstract, PostScript, DVI. Catarina Coquand. Introduction to ALF — a background. In 6th NWPT, page 157. vi+483 pp.

[Fal] Abstract, PostScript, DVI. Göran Falkman. Program separation as a basis for definitional higher order programming. In 6th NWPT, pages 158–172. vi+483 pp.

[Fre] Abstract, PostScript, DVI. Øyvind Bolme Fredriksen. General algebras. In 6th NWPT, pages 173– 187. vi+483 pp.

[HHK] Abstract, PostScript, DVI. Martin Hansen, Hans Hüttel, and Josva Kleist. Bisimulations for asynchronous mobile processes. In 6th NWPT, pages 188–189. vi+483 pp.

[HK] Abstract, PostScript, DVI. Claus Hintermeier and Hélène Kirchner. G^{2}-algebras. In 6th NWPT, pages 190–191. vi+483 pp.

[HP] Abstract, PostScript, DVI. Leszek Holenderski and Axel Poigné. Boolean automata: A compact representation of synchronous reactive systems. In 6th NWPT, pages 192–193. vi+483 pp.

[Har] Abstract, PostScript. Mait Harf. Structural synthesis of programs using regular data structures. In 6th NWPT, pages 194–202. vi+483 pp.

[Ing] Abstract, PostScript, DVI. Anna Ingólfsdóttir. A semantic theory for value-passing processes, late approach, part I: A denotational model and a complete axiomatization. In 6th NWPT, pages 203–219. vi+483 pp.

[Jen] Abstract, PostScript, DVI. Claus Torp Jensen. Interpreting broadcast communication in CCS with priority choice. In 6th NWPT, pages 220–236. vi+483 pp.

[Jeu] Abstract, PostScript, DVI. Johan Jeuring. Polytypic programming — abstract. In 6th NWPT, page 237. vi+483 pp.

[LSW] Abstract, PostScript, DVI. Kim G. Larsen, Bernhard Steffen, and Carsten Weise. A constraint oriented proof methodology based on modal transition systems. In 6th NWPT, pages 238–250. vi+483 pp.

[Las] Abstract, PostScript, DVI. Søren Bøgh Lassen. Reasoning with actions. In 6th NWPT, pages 251– 265. vi+483 pp.

[MD] Abstract, PostScript, DVI. Karoline Malmkjær and Olivier Danvy. Preprocessing by program specialization. In 6th NWPT, pages 266–268. vi+483 pp.

[Mag] Abstract, PostScript, DVI. Lena Magnusson. Introduction to ALF — an interactive proof editor. In 6th NWPT, page 269. vi+483 pp.

[NN] Abstract, PostScript, DVI. Hanne Riis Nielson and Flemming Nielson. Static and dynamic processor allocation for higher-order concurrent languages. In 6th NWPT, pages 270–285. vi+483 pp.

[Olv] Abstract, PostScript, DVI. Peter C. Ølveczky. Termination of order-sorted rewriting. In 6th NWPT, pages 286–299. vi+483 pp.

[PO] Abstract, PostScript, DVI. Jens Palsberg and Patrick O’Keefe. A type system equivalent to flow analysis. In 6th NWPT, pages 300–316. Extended abstract of a paper in Proc. POPL ’95, 22nd Annual SIGPLAN–SIGACT Symposium on Principles of Programming Languages.

[PPS] Abstract, PostScript, DVI. Wiesław Pawłowski, Paweł Pa̧czkowski, and Stefan Sokołowski. Specifying and verifying parametric processes. In 6th NWPT, pages 317–331. vi+483 pp.

[RSZ] Abstract, PostScript, DVI. Rimvydas Rukšėnas, Kaisa Sere, and Yi Zhao. On the formal derivation of a FEAL microprocessor. In 6th NWPT, pages 332–345. vi+483 pp.

[SNN] Abstract, PostScript, DVI. Kirsten Lackner Solberg, Hanne Riis Nielson, and Flemming Nielson. Strictness and totality analysis. In 6th NWPT, page 346. vi+483 pp.

[SW] Abstract, PostScript, DVI. K. Sere and Marina Waldén. Backward refinement for verifying distributed algorithms. In 6th NWPT, page 347. vi+483 pp.

[Sak] Abstract, PostScript, DVI. Jūratė Sakalauskaitė. Nonclausal resolution system for branching temporal logic. In 6th NWPT, pages 348–358. vi+483 pp.

[Sana] Abstract, PostScript, DVI. David Sands. Composed reduction systems. In 6th NWPT, pages 359– 376. vi+483 pp.

[Sanb] Abstract, PostScript, DVI.

David Sands. Towards operational semantics of contexts in functional languages. In 6th NWPT, pages 377–384. vi+483 pp.

[Sas] Abstract, PostScript, DVI.

Vladimiro Sassone. An approach to the category of net computations. In 6th NWPT, pages 385–399. To appear in InProceedings of TAPSOFT ’95.

[Tor] Abstract, PostScript, DVI.

Olof Torgersson. Functional logic programming in GCLA. In 6th NWPT, pages 400–414. vi+483 pp.

[Uus] Abstract, PostScript, DVI.

Tarmo Uustalu. Extensions of structural synthesis of programs. In 6th NWPT, pages 415–427. vi+483 pp.

[Wei] Abstract, PostScript, DVI.

Martin Weichert. Algebra of broadcasting systems: Value passing, sequential composition, and fork. In 6th NWPT, pages 428–443. vi+483 pp.

[vB] Abstract, PostScript, DVI.

Franck van Breugel. From branching to linear metric domains (and back). In 6th NWPT, pages 444–447. Abstract.

[AKLN] Abstract, PostScript.

Jørgen H. Andersen, Kåre J. Kristoffersen, Kim G. Larsen, and Jesper Niedermann. Automatic synthesis of real time systems. In 6th NWPT, pages 448–464. vi+483 pp.

[CM] Abstract, PostScript, DVI.

Catarina Coquand and Lena Magnusson. Demonstration of ALF. In 6th NWPT, page 465. vi+483 pp.

[SFC+] Abstract, PostScript.

Bernhard Steffen, B. Freitag, A. Claßen, Tiziana Margaria, and U. Zukowski. Intelligent software synthesis in the DaCapo environment. In 6th NWPT, pages 466–481. vi+483 pp.

[Sor] Abstract, PostScript, DVI.

Ib Holm Sørensen. Demonstration of the B-Toolkit. In 6th NWPT, pages 482–483. vi+483 pp.

[NS-94-6] Uffe H. Engberg, Kim G. Larsen, and Peter D. Mosses, editors. Proceedings of the 6th Nordic Workshop on Programming Theory (Aarhus, Denmark, 17–19 October, 1994), December 1994. vi+483 pp.